# Robust High-Fidelity Quantum Entanglement Distribution over Large-Scale Metropolitan Fiber Networks with Co-propagating Classical Signals

### v1

@@ Abstract
The Quantum Internet, a network of quantum-enabled infrastructure, represents the next frontier in telecommunications, promising unprecedented capabilities that cannot be attained by classical counterparts. A crucial step in realizing such large-scale quantum networks is the integration of entanglement distribution within existing telecommunication infrastructure. Here, we demonstrate a real-world scalable quantum networking testbed deployed within Deutsche Telekom’s metropolitan fibers in Berlin. Using commercially available quantum devices and standard add-drop multiplexing hardware, we distribute polarization-entangled photon pairs over dynamically selectable fiber paths ranging from 10 m to 82 km. Quantum signals, transmitted at 1324 nm (O-band), coexist with conventional bidirectional C-band traffic without dedicated fibers or infrastructure changes. Active stabilization of the polarization enables robust long-term performance, achieving entanglement fidelities between 85–99% and Clauser-Horne-Shimony-Holt parameter $S$-values between 2.36–2.74 during continuous multiday operation. By achieving high-fidelity entanglement distribution under real urban conditions with less than 1% downtime, we confirm the feasibility of hybrid quantum-classical networks under real-world conditions at metropolitan scale. These results establish deployment benchmarks and provide a practical roadmap for telecom operators to integrate quantum capabilities.
@@ I Introduction
The rapidly advancing field of quantum technology promises revolutionary applications ranging from quantum computing [1, 2, 3] to enhanced metrology [4, 5, 6, 7] and secure communications [8, 9, 10, 11, 12, 13], which will impact diverse areas such as financial transactions [14] and voting systems [15]. At the core of many of these applications lies entanglement, a distinctive quantum correlation that can be spatially separated, providing the foundation for quantum networks often referred to as the Quantum Internet. Entanglement-based networks underpin key future capabilities such as quantum repeaters [16], distributed sensing [7], quantum computing [17], and time synchronization [18].
@@
Photons, ideal carriers for this quantum information due to their weak environmental coupling, will require distribution through existing fiber optics to enable future quantum technologies. However, distributing entangled states over fiber-based telecom infrastructure that already transmits classical data is nontrivial. Key technical hurdles include the development of field-deployable quantum hardware, compatibility with legacy systems, and mitigation of noise, environmental fluctuations, and channel crosstalk. Prior demonstrations of quantum-classical coexistence have mostly been limited to single-channel [19, 20, 21] or controlled laboratory settings [22, 23, 24], with minor emphasis on long-term stability or scalability in real-world or multi-user metropolitan networks. Recently, more realistic approaches are starting to be explored [25, 26, 12].
@@
Figure 1: The BearlinQ layout. Setup schematic for adaptive routing of polarization-entangled photon pairs between local (T-labs) and deployed fiber paths, shown in the lower left with a schematic of DT’s Berlin fiber network nodes from the Winterfeldtstrasse laboratories (WFD). Fibers are used individually or stitched to reach $\sim 15$ km (HHI), $\sim 30$ km (Dotti), $\sim 60$ km (2$\times$ Dotti), or $\sim 82$ km (Dotti & HUB). Polarization-entangled photons are generated at the SRC device optically driven by the LSR. The idler photon ($795,\text{nm}$) is routed to the measurement station, while the signal photon ($1324,\text{nm}$) travels through the APC-INJ/CMP stack, fiber-gapped by the intervening path-length segments, with the path selected by an optical switch. OADMs after the APC-INJ and before the APC-CMP allow co/counter-propagation of C-band channels. After the APC-CMP, signal photons go to the measurement station. A high-pass filter suppresses classical leakage at the SNSPD. Coincidences at given polarizations are logged using a Swabian Instruments TDC, while CHSH measurements are enabled by rotating $\lambda/2$ plates and polarizing beam splitters (PBS).
@@
Quantum network deployments hinge on reliable instrumentation to generate, transmit, and measure entanglement as well as on systems to stabilize the overall infrastructure. A key design choice is the photonic degree of freedom, the most popular examples being time-bins [27], frequency-bins [13], or polarization [28, 29]. Among these, polarization encoding is particularly appealing because of its natural compatibility with other quantum systems such as quantum memories. However, photon polarization is highly sensitive to birefringence fluctuations which degrade entanglement fidelity and protocol reliability, an issue magnified in multi-path topologies common to real-world networks. This makes active polarization compensation essential for stable entanglement distribution.
@@
We demonstrate the integration of stable, high-fidelity polarization entanglement into an existing classical network under the city of Berlin. We call this hybrid quantum-classical network installation BearlinQ, a name reflecting our prototype’s technical robustness while evoking the strength and resilience symbolized by Berlin’s iconic bear. This is attained by distributing polarization-entangled O-band (1324 nm) photons over Deutsche Telekom’s (DT) metropolitan fiber network in the presence of co- and counter-propagating C-band (1550 nm) classical traffic under real-world conditions. Robust quantum performance is enabled by Qunnect’s integrated hardware suite comprised of an entanglement source, pumping lasers, and automatic polarization compensation to dynamically stabilize polarization across deployed fiber links. Integration with existing network infrastructure is made possible by optical add-drop multiplexers (OADMs), dynamic path switching, and adherence to International Telecommunication Union - Telecommunication Standardization Sector (ITU-T) channel standards, all without the need for dedicated fibers. To demonstrate network flexibility, we perform drift and switching experiments on different fiber paths. Long-term stability is tested for multiple days by alternating between local and deployed fiber segments of up to $\sim$82 km. Clauser-Horne-Shimony-Holt (CHSH) inequality [30] and fidelity measurements are periodically performed, confirming strong, continued quantum correlations in all scenarios. With automatic compensation, we maintain high fidelity with long-term quantum stability, demonstrating that reliable and scalable quantum networking is achievable by operators such as DT on existing telecom infrastructure, bringing quantum networking closer to real-world deployment.
@@
The paper is organized into two main sections outlining the network implementation and performance. Section II describes the BearlinQ fiber testbed, covering key hardware elements such as the entanglement source, polarization stabilization, and quantum-classical signal integration. Section III presents experimental results, including stability tests over urban distances, path-switching effects, and key quantum operational metrics for network operators.
@@ II.1 The BearlinQ network
Our quantum-classical hybrid network is centered around an entanglement generation hub connected to deployed fibers. A schematic of the layout is shown in Figure 1. Entanglement is generated by an off-the-shelf commercial device (SRC, see Section II.2) producing entangled photon pairs at 795 nm (NIR, idler) and 1324 nm (telecom O-band, signal), driven by a dedicated laser system (LSR). These rack-mounted (2U) devices enable deployment and comply with telecom collocation standards. The 795 nm photons are routed through local fibers directly to a measurement station. The 1324 nm photons are directed to the same measurement station either along a short local fiber or through deployed fibers that exit the laboratory via a series of patch panels, traverse the BearlinQ network, and loop back to the lab. A 2$\times$2 mechanical optical cross-switch allows dynamic selection of the fiber paths. Classical communication signals required for coordination are transmitted over the C-band and combined with/separated from the quantum channel using OADMs. A central control server orchestrates the experiment and analysis.
@@
To compensate for polarization drifts in the fibers caused by environmental fluctuations or switching events, an automatic polarization compensator (APC) actively maintains the polarization state of the O-band photons. Following entanglement generation, the quantum signal is sent through the device pair, which interweaves classical probing light into the fiber to enable multi-state polarization tracking and applies dynamic corrections to mitigate polarization drifts by time-multiplexing to ensure that any arbitrary polarization will be preserved during long-haul propagation (see Section II.3).
@@
At the measurement station, O-band and NIR photons are analyzed. Detection is performed using superconducting nanowire single-photon detectors (SNSPDs). Arrival times are time-stamped with a time-to-digital converter (TDC) and processed at the server to extract the CHSH $S$-value. The measurement station, consisting of motorized half-wave ($\lambda/2$) plates, fixed polarizing beam splitters, and a bandpass filter to suppress C-band noise, supports automated CHSH testing. Manual controllers fine-tune the local polarization reference and ensure alignment with the SNSPDs. This fine-tuning is required after any setup modification. Run-to-run variations in the measured $S$-values are primarily attributed to uncertainties in this manual calibration procedure (see Supp. Mat. IV.1).
Finally, we characterize our network by performing optical time domain reflectometer (OTDR) measurements on each deployed length of fiber. Explicit characterization is essential as real-world fibers are subject to a range of imperfections including environmental influences, splices, connector interfaces, and inconsistencies in manufacturing or installation which can introduce losses exceeding nominal specifications. This is evident in BearlinQ, where the 30 km and 60 km loops exhibit losses close to those expected, but the 15 km and 82 km loops contain segments with larger attenuation (see Figure 2).
@@
Figure 2: Fiber characterization. OTDR measurements for selected fibers of $\sim$ 15, 30, 60, and 82 km. Trace discontinuities correspond to fiber splices, connectors, and mating sleeves. For reference, the expected attenuation of standard single-mode fiber (SMF-28) at 1324 nm is $\sim 0.32$ dB/km [31], as indicated by the black dashed line.
@@ II.2 Quantum-Classical Integration and Entanglement Source
Important considerations for the design of hybrid networks that need to balance the requirements of both quantum operations and everyday classical use are the choice of relative wavelengths for the classical and quantum signals and the intimately-related selection of an entanglement source. The former is mainly informed by two elements: channel density and noise management. Transmitting quantum and classical signals respectively in the O- and C-bands effectively addresses both issues. Additionally, the selected O-band entanglement source meets necessary reliability benchmarks for stable real-life operation.
The high channel density of the C-band supports a large number of simultaneous users through a mature and standardized dense wavelength division multiplexing (DWDM) infrastructure with narrow channel spacings of 50 GHz. In contrast, the O-band is limited to coarse WDM (CWDM) standards, typically offering only three usable channels with 20 nm spacing, which substantially limits its multiplexing capacity. This disparity naturally assigns classical signals to the C-band, where high data throughput and multi-service access are essential for existing network users, while reserving the O-band for low-noise quantum channels.
Single-photon signals are extremely susceptible to noise in the form of uncorrelated photons that make their way into the measurement devices. The dominant noise source is spontaneous Raman scattering [32], which occurs when light interacts with its propagation medium generating broadband frequency-shifted photons that overlap with the quantum signal and thus cannot be filtered. The accurate selection of the relative wavelengths of the quantum and classical channels effectively mitigates this issue. Raman scattering is an asymmetric process, predominantly producing Stokes-shifted (longer wavelength) photons. Transmitting quantum signals in the O-band and classical data in the C-band is particularly favorable. By placing classical channels at longer wavelengths, the majority of Raman noise falls outside the quantum detection window, significantly avoiding in-band noise pollution. This approach enables the use of high classical launch powers (¿10 dBm), relaxes spectral filtering requirements (reducing insertion loss), and enhances system robustness for real-world deployment of coexisting channels over existing fiber infrastructure.
Having identified the O-band as the ideal range for quantum signals, we employ Qunnect’s SRC as the source of entangled photons. The SRC operates at room-temperature delivering ¿$10^{7}$ photon pairs/s with $\sim 1$ GHz linewidth (high brightness) and intrinsically stable polarization entanglement with large signal-idler cross-correlations ($g^{(2)}_{si}\gtrsim 30$) [33] and a heralding efficiency of $\sim 30\%$ [28]. Its O-band wavelength also leverages the fibers’ 1310 nm low-dispersion window and removes the need for frequency conversion. These features make the SRC an ideal tool for the generation of entanglement in quantum networks run by telecom operators, as it eliminates many constraints of conventional sources such as limited brightness or fidelity, or the need for cryogenics, vacuum apparatus, or controlled laboratory environments.
@@
The photon pair generation rate scales linearly with pump laser power, as the probability of pair creation is directly proportional to the number of pump photons available for the nonlinear interaction [28]. Increasing the pump power is particularly useful in compensating for losses that accumulate with increasing fiber length, as a higher pair-generation rate helps maintain sufficient photon rates for use-cases. However, while increasing pump power helps overcome larger transmission losses, it also introduces a trade-off in signal-idler correlation strength, as quantified by the cross-correlation function $g^{(2)}_{si}$ [33]. The inverse relationship between pair-rate and $g^{(2)}_{si}$ imposes a practical limitation, as large pump powers ultimately reduce the proportion of usable entangled pairs. To accommodate different fiber lengths while ensuring high fidelity a constant coupling power of $8.4$ mW was used, along with variable pump powers of $\sim 450$ µW and $\sim 900$ µW.
@@ II.3 Active Polarization Stabilization
Environmental factors such as temperature fluctuations [34] and mechanical stress [35] on single-mode, non-polarization maintaining optical fibers induce birefringence changes, causing unwanted rotations of the state of polarization of propagating photons. Protocols relying on polarization require precise control of these states to correctly analyze quantum correlations. Uncompensated variations can effectively randomize the polarization basis, degrading the fidelity and usability of entanglement across the network. To ensure accurate correlation measurements, we actively track and compensate for polarization drifts by maintaining a stable and distributed polarization reference across all nodes. The APC system (injector, APC-INJ, and compensator, APC-CMP) provides the reference and stabilizes the fibers to it. The APC-INJ transmits a known control sequence of approximately the same wavelength as the qubits to the APC-CMP at the opposite end of a fiber link (see Supp. Mat. IV.1). The APC-CMP then analyzes the deviations of a sequence of states from their intended polarizations and applies corrections through a low-loss elasto-electro-optical feedback mechanism [29]. This system uses time-multiplexing to operate over the same fibers as both quantum and classical signals.
Figure 3: Passive fiber operation. Time evolution of $S$-value and quantum fidelity, $F_{Q}$, for a local (blue) and a 60 km deployed (magenta) fiber path, both without active stabilization control. The red dashed line indicates the CHSH threshold at $S=2$, the classical limit for local hidden variable theories. A value $S>2$ is required to certify entanglement and enable quantum applications. A high quantum fidelity (high similarity of the transmitted state to a well-defined reference state) is also necessary for channel suitability, but is not guaranteed by $S>2$. The local fibers within the T-Labs environment exhibit a very stable $S$ and $F_{Q}$ over $\sim 60$ hours. In contrast, the outside path exhibits a sudden drop in both below usable thresholds. Error bars represent $1\sigma$ uncertainties from Poissonian count statistics propagated through CHSH calculations. The integration time for each waveplate configuration was 1 s (2 s) for the local (60 km) path.
@@
We characterize the fibers to assess drift rate and evaluate the required period of basis re-alignments by performing CHSH and quantum fidelity, $F_{Q}$, measurements over both local and deployed fibers in the absence of active stabilization. The CHSH $S$-value is calculated using standard methods [36], while the upper/lower quantum fidelity bound calculation is detailed in Supp. Mat. IV.2. As shown in Figure 3, the fiber segments flanking the APC pair (common to both the local and deployed paths) and the local T-Labs fiber linking the two APCs exhibit high stability over multiple days. In contrast, while the laboratory environment effectively suppresses drifts, the deployed loop can be subject to instability and unpredictability. In the experimental run presented here, the selected fibers demonstrate resilience to slow drifts but exhibit sudden jumps. This behavior is not consistent across deployments: in other instances, gradual drifts are more prevalent [37]. This variation likely arises from differences in installation practices: specifically, here all fibers are buried, and the infrastructure is modern and actively maintained by DT. Nevertheless, we confirm that active stabilization is required for stable network operation by telecom operators.
Refer to captionRefer to caption
Figure 5: Sustained automated multi-path switching in a metropolitan network. a) Sequence of logic operations showing key timing intervals for CHSH measurement (blue, $t_{\mathrm{CHSH}}$), active polarization control (green, $t_{\mathrm{APC\,comp}}$), path switching (yellow, $T_{\mathrm{path\,switch}}$), and local acquisition (yellow, $t_{\mathrm{local}}$). Vertical dotted lines mark switching events. b) Start and end classical fidelity for all APC routines during the switch from BearlinQ to local fiber. Histograms at right show the distribution of final classical fidelities measured by the APC. c) Local long-term CHSH and quantum fidelity traces, each corresponding to a switch from one of the deployed fibers (see following panels); similar color tones indicate the same measurement run. d), e), f) Long-term CHSH and quantum fidelity over 15 km, 30 km, and 60 km, respectively. Error bars in c–f show $1\sigma$ uncertainties from Poissonian count statistics propagated through the CHSH $S$ calculation. In c–f, the red dashed line marks $S=2$. The integration time for each waveplate configuration in the CHSH and fidelity measurements was 1 s for the local, 15 km, and 30 km paths, and 2 s for the 60 km path.

### v2

# High-Fidelity Quantum Entanglement Distribution in Metropolitan Fiber Networks with Co-propagating Classical Traffic

@@ Abstract
The Quantum Internet, a network of quantum-enabled infrastructure, represents the next frontier in telecommunications, promising capabilities that cannot be attained by classical counterparts. A crucial step in realizing such large-scale quantum networks is the integration of entanglement distribution within existing telecommunication infrastructure. Here, we demonstrate a real-world scalable quantum networking testbed deployed within Deutsche Telekom’s metropolitan fibers in Berlin. Using commercially available quantum devices and standard add-drop multiplexing hardware, we distributed polarization-entangled photon pairs over dynamically selectable fiber paths ranging from 10 m to 60 km, and showed entanglement distribution over up to approximately 100 km. Quantum signals, transmitted at 1324 nm (O-band), coexist with conventional bidirectional C-band traffic without dedicated fibers or infrastructure changes. Active stabilization of the polarization enables robust long-term performance, achieving entanglement Bell-state fidelity bounds between 85–99% and Clauser-Horne-Shimony-Holt parameter $S$-values between 2.36-2.74 during continuous multiday operation. By achieving a high-fidelity entanglement distribution with less than 1.5% downtime, we confirm the feasibility of hybrid quantum-classical networks under real-world conditions at the metropolitan scale. These results establish deployment benchmarks and provide a practical roadmap for telecom operators to integrate quantum capabilities.
@@ I Introduction
The field of quantum technology promises revolutionary applications ranging from quantum computing [1, 2, 3, 4] to enhanced metrology [5, 6, 7, 8] and secure communications [9, 10, 11, 12, 13, 14], which will impact diverse areas such as financial transactions [15] and voting systems [16]. At the core of many of these applications lies entanglement, a distinctive quantum correlation that can be spatially separated, providing the foundation for quantum networks often referred to as the Quantum Internet. Entanglement-based networks underpin key future capabilities such as quantum repeaters [17], distributed quantum sensing [8], quantum computing [18], and time synchronization [19].
@@
Photons, ideal quantum information carriers because of weak environmental coupling, often require or benefit from distribution through existing fiber optics to enable quantum applications. However, transmitting entangled states over a fiber infrastructure that already transmits classical data is a challenge. Key technical hurdles include the development of reliable and field-deployable quantum hardware, compatibility with legacy systems, and mitigation of noise, environmental fluctuations, and channel crosstalk. Previous demonstrations of quantum-classical coexistence have mostly been limited to single-channel [20, 21, 22] or controlled laboratory settings [23, 24, 25], with minor emphasis on long-term stability and multi-user metropolitan networks. Recently, more realistic approaches are starting to be explored [26, 27, 13].
@@
Figure 1: The BearlinQ layout. Diagram of the setup for adaptive routing of polarization-entangled photon pairs between local (T-labs) and deployed fiber paths, shown in the lower left with a schematic of DT’s Berlin fiber network nodes from the Winterfeldtstrasse laboratories (WFD). Polarization-entangled photons are generated at the SRC device optically driven by the LSR. The idler photon (795 nm) is routed to the measurement station, while the signal photon (1324 nm) travels through the APC-INJ/CMP stack, fiber-gapped by the intervening path-length segments, with the path selected by an optical switch. OADMs after the APC-INJ and before the APC-CMP allow co/counter-propagation of C-band channels. After the APC-CMP, signal photons go to the measurement station. Coincidences at given polarizations are logged using a Swabian Instruments TDC, while CHSH measurements are enabled by rotating $\lambda/2$ plates and polarizing beam splitters (PBS).
@@
The photonic degree of freedom used to encode entanglement is a key network design choice. The most popular are time-bin [28], frequency-bin [14], and polarization [29, 30]. Among these, polarization encoding is particularly appealing because of its ease of manipulation and its natural compatibility with other quantum systems. However, photon polarization is highly sensitive to birefringence fluctuations, which degrade entanglement fidelity and protocol reliability, an issue magnified in multipath topologies common to real-world networks. This makes active polarization compensation essential for stable entanglement distribution.
@@
Here, we demonstrate the integration of high-fidelity polarization entanglement into an existing classical network under the city of Berlin. We call this hybrid quantum-classical network BearlinQ, a name reflecting the technical robustness of our prototype while evoking the strength and resilience symbolized by Berlin’s iconic bear. The network installation consists of distributing polarization-entangled O-band (1324 nm) photons over Deutsche Telekom’s (DT) metropolitan fiber network in the presence of co- and counter-propagating C-band (1550 nm) classical traffic under real-world conditions. Robust quantum performance is enabled by Qunnect’s integrated hardware suite comprised of an entanglement source, pumping lasers, and automatic polarization compensation to dynamically stabilize polarization across deployed fiber links. Integration with existing network infrastructure is made possible by optical add-drop multiplexers (OADMs), dynamic path switching, and adherence to International Telecommunication Union - Telecommunication Standardization Sector (ITU-T) channel standards, all without the need for dedicated fibers. To demonstrate network flexibility, we performed drift and switching experiments on different fiber paths. Long-term stability is confirmed by running the experiment over multiple days by alternating between local and deployed fiber segments of up to $\sim$60 km. Clauser-Horne-Shimony-Holt (CHSH) inequality [31] and fidelity measurements are periodically performed, confirming strong, continued quantum correlations in all scenarios. With automatic compensation, we maintain high fidelity with long-term quantum stability, demonstrating that operators can achieve reliable and scalable quantum networking on existing telecom infrastructure.
@@
The paper is organized into two main sections that outline the network implementation and performance. Section II describes the BearlinQ fiber testbed, covering key hardware elements such as the entanglement source, polarization stabilization, and quantum-classical signal integration. Section III presents experimental results, including stability tests over urban distances, path-switching effects, and key metrics for network operators.
@@ II.1 The BearlinQ network
Our quantum-classical hybrid network is centered around an entanglement generation hub connected to deployed fibers. A schematic of the layout is shown in Figure 1. Entanglement is generated by an off-the-shelf commercial device (SRC, see Section II.2) producing entangled photon pairs at 795 nm (NIR, idler) and 1324 nm (telecom O-band, signal), driven by a dedicated laser system (LSR). These rack-mounted (2U) devices enable deployment and comply with telecom collocation facility requirements. The 795 nm photons are routed through local fibers directly to a measurement station. The 1324 nm photons are directed to the same measurement station either along a short local fiber or through long-haul deployed fibers that exit the laboratory via a series of patch panels, traverse the BearlinQ network, and loop back to the lab. A MEMs fiber optical cross-switch allows dynamic selection of fiber paths. Classical communication signals required for coordination are transmitted over the C-band and combined with/separated from the quantum channel using OADMs. A central control server orchestrates the experiment and analysis.
@@
To compensate for polarization drifts in the fibers caused by environmental fluctuations or switching events, the polarization state of the O-band photons is actively maintained by an automatic polarization compensator (APC). The APC consists in a pair of devices which, following entanglement generation, accept the quantum signal and interweave classical probing light at the same wavelength into the network fibers. This time-multiplexed classical signal enables multi-state polarization tracking and the application of dynamic corrections to mitigate polarization drifts and thus ensure that any arbitrary polarization will be well preserved during long-haul propagation (see Section II.3).
@@
At the measurement station, O-band and NIR photons are analyzed. Detection is performed using superconducting nanowire single-photon detectors (SNSPDs). Arrival times are time-stamped with a time-to-digital converter (TDC) and processed at the server to extract the CHSH $S$-value. The measurement station, consisting of motorized half-wave ($\lambda/2$) plates, fixed polarizing beam splitters, and a bandpass filter to suppress C-band noise, supports automated CHSH tests. Manual controllers calibrate the local polarization reference and also ensure alignment with the polarization sensitive SNSPDs. Run-to-run variations in the measured $S$-values are primarily attributed to uncertainties in this manual calibration procedure (see Supp. Mat. IV.1).
@@
Figure 2: Fiber characterization. OTDR measurements of fibers used individually or stitched together to reach 15$\sim$ 15∼ 15 km (HHI), 30$\sim$ 30∼ 30 km (Dotti), 60$\sim$ 60∼ 60 km (2$\times$ Dotti), or 100$\sim$ 100∼ 100 km (Strausberg) (specifically: 13.6, 30.5, 61.0 and 97.9 km). Trace discontinuities correspond to splices, connectors, and mating sleeves. For reference, the expected attenuation of standard SMF-28 single-mode fiber at 1310 nm is 0.31$\sim$ 0.31∼ 0.31 dB/km, as indicated by the dashed black line.
Finally, we characterize the network by performing optical time-domain reflectometer (OTDR) measurements on each length of fiber. Explicit characterization is essential because real-world fibers are subject to a range of imperfections including environmental influences, splices, connector interfaces, and inconsistencies in manufacturing or installation that can introduce losses exceeding nominal specifications. This is evident in BearlinQ, where the 30 km and 60 km paths exhibit losses close to those expected, but the 15 km and 100 km stretches contain segments with greater attenuation (see Figure 2).
@@ II.2 Quantum-Classical Integration and Entanglement Source
A key consideration in the design of networks that support both quantum operations and conventional classical traffic is the selection of relative wavelengths for the two domains, along with the closely related choice of entanglement source. This decision is primarily informed by two factors: spectral channel density and noise management. Assigning classical signals to the C-band and quantum signals to the O-band effectively addresses both concerns. The C-band’s high channel density enables support for a large number of simultaneous users through a mature and standardized dense wavelength division multiplexing (DWDM) infrastructure, featuring narrow 50GHz channel spacing. In contrast, the O-band is limited to coarse WDM (CWDM) standards, typically offering only three usable channels spaced by 20 nm, which significantly constrains its multiplexing capacity. From a telecom operator’s perspective, this disparity naturally favors allocating classical signals to the C-band, which is better suited to high user density, bandwidth-intensive applications, and diverse service requirements. Conversely, the O-band is better suited for low-noise quantum channels, which are more appropriate for transmitting fragile entangled states.
Single-photon signals are extremely susceptible to noise in the form of uncorrelated photons that make their way into the measurement devices. Once polarization is stabilized through the network, the relative rate of the quantum signal and of the noise background is the primary limiting factor to the quality of the measured entanglement as the distances increase. The dominant inter-band noise source in optical fibers is spontaneous Raman scattering [32], which occurs when light interacts with its propagation medium generating broadband frequency-shifted photons that overlap with the quantum signal and thus cannot be filtered. Proper selection of the relative wavelengths of the quantum and classical channels can effectively mitigate this issue. Raman scattering is an asymmetric process, predominantly producing Stokes-shifted (longer wavelength) photons. Transmitting quantum in the O-band and classical data in the C-band is thus particularly favorable. By placing classical channels at longer wavelengths, the majority of Raman noise falls outside the quantum detection window, significantly avoiding in-band noise pollution. This approach enables the use of high classical launch powers, relaxes spectral filtering requirements (reducing insertion loss) and enhances the robustness of the system for deployment of coexisting channels over existing fiber infrastructure.
[…]
@@
The photon pair generation rate scales linearly with pump laser power, as the probability of pair creation is directly proportional to the number of pump photons available for the nonlinear interaction [29]. Increasing the pump power is particularly useful in compensating for losses that accumulate with increasing fiber length, as a higher pair-generation rate helps maintain sufficient photon rates for use-cases. However, while increasing pump power helps overcome larger losses, it also introduces a trade-off in signal-idler correlation strength, as quantified by the cross-correlation function $g^{(2)}_{si}$ [33]. The inverse relationship between pair-rate and $g^{(2)}_{si}$ imposes a practical limitation, as large pump powers ultimately reduce the proportion of usable entangled pairs. To accommodate different fiber lengths while ensuring high-fidelity a constant coupling power of 8.48.48.4 mW was used, along with variable pump powers of 450900$\sim$ 450-900∼ 450 - 900 µW.
@@ II.3 Active Polarization Stabilization
Environmental factors such as temperature fluctuations [34] and mechanical stress [35] on single-mode, non-polarization maintaining optical fibers induce birefringence changes, causing unwanted rotations of the state of polarization of propagating photons. Protocols relying on polarization require precise control of these states to correctly analyze correlations. Uncompensated variations can effectively randomize the basis, degrading the fidelity and usability of entanglement across the network.
To ensure accurate correlation measurements, we actively track and compensate for polarization drifts by maintaining a stable and distributed polarization reference across all nodes. The APC system (injector, APC-INJ, and compensator, APC-CMP) provides such a reference and stabilizes the fibers to it. The APC-INJ transmits a known control sequence of approximately the same wavelength as the qubits, time-multiplexed with the quantum channel, to the APC-CMP at the opposite end of a fiber link (see Supp. Mat. IV.1). The APC-CMP analyzes the deviations of a sequence of states from their intended polarizations and applies corrections through a low-loss and rapid feedback mechanism [30].
Figure 3: Passive operation. Time evolution representative example of the CHSH $S$-value for both the local path (green) and a 30 km deployed fiber (purple), measured without active polarization stabilization [36]. The common fiber segments flanking the APC pair, as well as the local short fiber, exhibit exemplary high stability. In contrast, the deployed path shows a gradual decline in $S$, with occasional abrupt drops, highlighting the dominant impact of environmental perturbations on long-term stability. The red line marks the threshold at $S=2$: the classical bound for local hidden variable theories. $S>2$ indicates entanglement, but does not guarantee high quantum channel fidelity. Error bars denote $1\sigma$ uncertainties derived from Poissonian propagation.
[…]
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
We characterize the fibers to assess the drift rate and evaluate the required period of basis realignments by performing CHSH measurements and quantum fidelity, $F_{Q}$, over both local and deployed fibers in the absence of active stabilization. The $S$-value is calculated using standard methods [37], while the upper/lower quantum fidelity bound calculation is detailed in Supp. Mat. IV.2. Figure 3 illustrates the necessity of automatic polarization stabilization in field-deployed optical fiber networks and establishes the fundamental premise for the studies presented in this work. It shows the results of a long-term entanglement distribution experiment presented in [36], which used a setup similar to that proposed in this work, but without active stabilization. As observed, the temporal evolution of the $S$-value and $F_{Q}$ for the selected fiber exhibits a gradual drift over time, punctuated by sporadic sudden jumps. This behavior is consistently observed across different deployments and is likely attributable to common installation practices. In this particular case, all fibers are buried, and the infrastructure is modern and actively maintained by DT. Nonetheless, our findings confirm that active polarization stabilization is essential to ensure continuous and stable network operation for telecommunication providers.